OpenStudy (anonymous):

Use Taylors Inequality to estimate the accuracy of the approximation f(x)≈Tn(x) 0.8≤x≤1.2 \[sin(1)+cos(1)(x-1)-\frac{sin(1)(x-1)^2}{2}-\frac{cos(1)(x-1)^3}{6}\] when x lies in the given interval

OpenStudy (anonymous):

so far we know that \[\left|R_n(x) \right|=\left|f(x)-T_n(x)\right|\]

OpenStudy (anonymous):

and we also know that if \[\left|f^{(n+1)}(x)\right|\le M\] then \[\left|R_n(x) \right|\le \frac{M}{(n+1)!}\left| x-a\right|^{n+1}\]

OpenStudy (anonymous):

is the first term is sin1 or sin t?

OpenStudy (anonymous):

the original equation is f(x)= sinx a=1 and n=3

OpenStudy (anonymous):

oh I thought wrong that were coff !!

OpenStudy (anonymous):